# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ 4.6 Bridges between trajectories in $\mathcal{N}$=4 SYM
In the Figure, we see that following the Bridge leads to a subleading Regge trajectory with twist-$6$, i.e. shifted by 4 just like in ABJM theory.
Finally, just like in ABJM theory the Bridge itself can be seen as spin-shadow of a further trajectory with twist-$4$. Thus, we have found a very nontrivial manifestation of the symmetry: namely the statement that the image of the Konishi operator can be found, at $S_{sl(2)}=0$ (i.e. below the unitarity bound), on a trajectory of twist-$4$. As already mentioned in the Introduction, this observation was independently made recently in Julius:2024ewf . The findings of the present paper indicate that the statement extends to all loops.
Figure 13: Here we show the existence of two continuous curves going through the Konishi operator (the red dot) belonging to the sl(2) sector. They are obtained with the two gluing matrices discussed in 3.3. The unitarity-allowed region is depicted in gray. Coupling $g=0.5$ is considered.
@@ 5 Discussion and Outlook
A surprise of our analysis was revealing a non-perturbative symmetry between non-BPS local operators multiplets and their spin-flips with $S_{\text{top}}\rightarrow d-2-S_{\text{top}}$ on the Chew-Frautschi plot [^30] This definition of the symmetry applies to the spin of the top component of the supermultiplet. The whole supermultiplet gets then carried along.. The symmetry operates formally through an interchange of twist and co-twist defined in the Introduction, and implies that, for every non-BPS local operator which living on a certain Regge trajectory, its spin shadow will be present on another Regge trajectory below the unitarity bound, i.e. in the region that is not described by the Lorentzian inversion formula. We stress that this symmetry is not explained trivially by supersymmetry, since it relates Regge trajectories that are not connected by the action of supercharges. Rather, it emerges through a nontrivial interplay of the form of the QSC equations (and the way they encode the Weyl symmetries), and the way their solutions behave. Since it is not possible to track rigorously the solutions of the nonlinear QSC equations, we do not have a complete mathematical proof of the symmetry, but we highlighted that it seems to work by a very simple and elegant mechanism: in particular, the heart of the matter is that every non-BPS local operator seems to lie at the intersection of two distinct continuous trajectories, one its own Regge trajectory and the other the spin flip of a separate subleading trajectory. The numerical study of several states has confirmed this property, which then implies the twist/co-twist symmetry.
@@
We also pointed out that exactly the same symmetry seems to be present in another example of integrable AdS/CFT, i.e. $\mathcal{N}$=4 SYM, and demonstrated it by a preliminary numerical exploration, where we highlighted that the spin shadow of the Konishi operator (which has twist $2$) lives on a subleading trajectory of twist-$4$. This is precisely what was recently observed at strong coupling in Julius:2024ewf . What seems to be the same symmetry was also observed at weak coupling in Henriksson:2017eej . Moreover, we notice that a slightly different symmetry, also related to Weyl reflections, was observed in non-planar, double trace data in Aprile:2017qoy . It would be very interesting to understand where the twist/co-twist symmetry comes from in ABJM and $\mathcal{N}$=4 SYM and possibly other holographic theories (it is natural to conjecture that a version of this symmetry would be present also in integrable AdS3/CFT2).
@@
Further, it would be nice to find an interpretation of this duality in the theory of light-ray operators, and to understand if there might be versions of the same symmetry in other CFTs. An interesting problem is also raised by the increase of degeneracies of trajectories with twist. What does this imply together with the symmetry? Two possible mechanisms likely both play a role: first, some higher-twist trajectories could have points below the unitarity bound that, despite having negative integer spin (in steps of 2), are not spin-flipped images of local operators, and, from our point of view, this would be highlighted by the fact that they satisfy QSC equations with non constant gluing matrix [^31] Coinciding with this, the analytic continuation of OPE coefficients would have zeros at these points Henriksson:2023cnh . In Julius:2024ewf , see Fig. 2, we see already some examples of this mechanism. Notice that the “extra” operators on subleading trajectories are not shadows of local operators, and therefore they are not involved in the twist/co-twist symmetry. ; in addition, some “extra” trajectories might bend before reaching these negative values of $S$. In general, it would be important to decode the full pattern in which trajectories are threaded together by the twist/co-twist symmetry.
@@
Finally, another motivation for developing tools for studying continuous spin is to gather a new type of data to use in the Bootstrability approach Cavaglia:2021bnz ; Cavaglia:2022qpg ; Cavaglia:2023mmu ; Caron-Huot:2022sdy . So far, this program has used the spectrum of low-lying local operators coming from QSC methods as an input for the conformal bootstrap in order to derive constraints on correlation functions. Since continuous spin comes with a tight structure in CFT Caron-Huot:2017vep , it would be very nice to explore if the input on Regge trajectories at any continuous $S$ coming from the QSC could be used to put even more constraints on OPE coefficients. This direction might be important especially to study Bootstrability for the full 3D or 4D theory (see Caron-Huot:2022sdy ), which is significantly more complex than the most studied defect Wilson line setup. In a parallel development, recently there was significant progress in the study of 4-point correlators in $\mathcal{N}$=4 SYM at strong coupling Alday:2022uxp ; Alday:2022xwz ; Alday:2023mvu ; Alday:2023jdk ; Alday:2023flc . The approach of these works uses the analytic conformal bootstrap, with input from integrability spectral data on the first Regge trajectories. This analysis has also led to new structural insight on the analytic structure of the spectrum itself Julius:2023hre ; Julius:2024ewf . The symmetry we are discussing might give further constraints and be helpful in this program.
@@ Acknowledgements
We are grateful to Mikhail Alfimov, Francesco Aprile, Till Bargheer, Carlos Bercini, Riccardo Conti, Nikolay Gromov, Tobias Hansen, Johan Henriksson, Julius Julius, Rob Klabbers, Shota Komatsu, Madalena Lemos, Marco Meineri, Michelangelo Preti, Nicolò Primi, Nika Sergeevna Sokolova, Istvan Szećsényi for discussions related to this work.
@@ Appendix A Weak coupling checks of the symmetry in ABJM theory
$\gamma_{2}$ $=4\left(S_{1}(j+1)+S_{-1}(j+1)\right)+\frac{2\left(1-(-1)^{1+j}\right)}{j+2}+8,$ $\text{for}\,\,\mathcal{V}^{4}_{3+2j,3},\;\;(\tau=2,\;j\in\mathbb{N}),$
$\gamma_{2}$ $=4\left(S_{1}(j+2)-S_{-1}(j+2)\right)+\frac{\left(1+(-1)^{j}\right)}{j+3}-\frac{\left(1-(-1)^{j}\right)}{j+2}+4,$ $\text{for}\,\,\mathcal{V}^{4}_{4+2j,4},\;\;(\tau=3,\;j\in\mathbb{N}),$
$\gamma_{2}$ $=4\left(S_{1}(j+2)+S_{-1}(j+2)\right)+\frac{2\left(1-(-1)^{j}\right)}{j+3}+\frac{32}{3},$ $\,\,\mathcal{V}^{4}_{5+2j,5},\;\;(\tau=4,\;j\in 2\mathbb{N}+1),$ (181)
@@
$j=\frac{\text{max}\left(\delta_{1},0\right)-\text{max}\left(\delta_{2},0\right)}{2},\;\;\;\Delta=L+\frac{\text{max}\left(\delta_{1},0\right)+\text{max}\left(\delta_{2},0\right)}{2}.$ (183)
@@
$\gamma_{2}=4\left(S_{1}(j+1)-S_{-1}(j+1)\right),\;\;\text{ for }\mathcal{V}^{2}_{3+2j},\;\;(\tau=1,\;j\in\mathbb{N}).$
@@
$\left.\mathcal{V}^{2}_{3+2j}\right|_{j\rightarrow 0},\;\;(\tau=1,\tilde{\tau}=2)\;\;\longleftrightarrow\;\;\left.\mathcal{V}^{4}_{3+2j,3}\right|_{j\rightarrow-1},\;\;(\tau=2,\tilde{\tau}=1).$ (184)
@@
$\left.\mathcal{V}^{4}_{3+2j,3}\right|_{j\rightarrow 0},\;\;(\tau=2,\tilde{\tau}=3)\;\;\longleftrightarrow\;\;\left.\mathcal{V}^{4}_{4+2j,4}\right|_{j\rightarrow-1},\;\;(\tau=2\tilde{\tau}=2),$ (185)
@@
$\left.\mathcal{V}^{4}_{4+2j,4}\right|_{j\rightarrow 0},\;\;(\tau=3,\tilde{\tau}=4)\;\;\longleftrightarrow\;\;\left.\mathcal{V}^{4}_{5+2j,5}\right|_{j\rightarrow-1},\;\;(\tau=4,\tilde{\tau}=3),$ (186)
@@
Notice that in general the symmetry connects a state of spin $j$ on a trajectory of twist $\tau$ to a point on a trajectory of twist $\tilde{\tau}=d-2+\tau+2j$, continued to unphysical spin $\tilde{j}=-d+2-j$. The only instances of the symmetry which we can check based on the analytic trajectories (A) are the ones we have presented, as we do not have the exact expressions in order to perform the analytic continuations to negative $j$ for higher twists.
@@ Appendix B Useful ancillaria
Here we report some additional internal constraints of the QSC construction.
The leading prefactors coefficients of the expansion (163) satisfy (see Bombardelli:2017vhk )

### v3

@@ 4.6 Bridges between trajectories in $\mathcal{N}$$=$$4$ SYM
In the Figure, we see that following the Bridge leads to a subleading Regge trajectory with twist $6$, i.e. shifted by 4 just like in ABJM theory.
Finally, just like in ABJM theory the Bridge itself can be seen as spin-shadow of a further trajectory with twist $4$. Thus, we have found a very nontrivial manifestation of the symmetry: namely the statement that the image of the Konishi operator can be found, at $S_{sl(2)}=0$ (i.e. below the unitarity bound), on a trajectory of twist-$4$. As already mentioned in the Introduction, this observation was independently made recently in Julius:2024ewf . The findings of the present paper indicate that the statement extends to all loops.
Figure 13: Here we show the existence of two continuous curves (i.e. a Regge trajectory and a Bridge) going through the Konishi operator, which is marked by the solid red bullet. They are obtained with the two gluing matrices discussed in 3.3. A portion of the unitarity-allowed region is depicted in gray: $11<S_{sl(2)}<\Delta_{sl(2)}+$ 1 Gromov:2014caa . The value of the coupling $h=0.5$ is considered.
@@ 5 Discussion and Outlook
A surprise of our analysis was revealing a non-perturbative symmetry between non-BPS local operators multiplets and their spin-flips with $S_{\text{top}}\rightarrow-d+2-S_{\text{top}}$ on the Chew-Frautschi plot [^34] This definition of the symmetry applies to the spin of the top component of the supermultiplet. The whole supermultiplet gets then carried along.. The symmetry operates formally through an interchange of twist and co-twist defined in the Introduction, and implies that, for every non-BPS local operator which living on a certain Regge trajectory, its spin shadow will be present on another Regge trajectory below the unitarity bound, i.e. in the region that is not described by the Lorentzian inversion formula. We stress that this symmetry is not explained trivially by supersymmetry, since it relates Regge trajectories that are not connected by the action of supercharges. Rather, it emerges through a nontrivial interplay of the form of the QSC equations (and the way they encode the Weyl symmetries), and the way their solutions behave. Since it is not possible to track rigorously the solutions of the nonlinear QSC equations, we do not have a complete mathematical proof of the symmetry, but we highlighted that it seems to work by a very simple and elegant mechanism: in particular, the heart of the matter is that every non-BPS local operator seems to lie at the intersection of two distinct continuous trajectories, one its own Regge trajectory and the other the spin flip of a separate subleading trajectory. The numerical study of several states has confirmed this property, which then implies the twist/co-twist symmetry.
@@
We also pointed out that exactly the same symmetry seems to be present in another example of integrable AdS/CFT, i.e. $\mathcal{N}$$=$$4$ SYM, and demonstrated it by a preliminary numerical exploration, where we highlighted that the spin shadow of the Konishi operator (which has twist $2$) lives on a subleading trajectory of twist-$4$. This is precisely what was recently observed at strong coupling in Julius:2024ewf . What seems to be the same symmetry was also observed at weak coupling in Henriksson:2017eej . Moreover, in these works the symmetry appears to hold for the full CFT data including OPE coefficients! We also notice that a slightly different symmetry, also related to Weyl reflections, was observed in non-planar, double trace data in Aprile:2017qoy . It would be very interesting to understand where the twist/co-twist symmetry comes from in ABJM and $\mathcal{N}$$=$$4$ SYM and possibly other holographic theories (it is natural to conjecture that a version of this symmetry would be present also in integrable AdS3/CFT2). One could also try to explore if something related happens in the Fishnet theory Gurdogan:2015csr ; Grabner:2017pgm ; Gromov:2017cja ; Levkovich-Maslyuk:2020rlp , where the $\mathcal{N}$$=$$4$ supersymmetry but also unitarity are broken.
Further, it would be nice to find an interpretation of this duality in the theory of light-ray operators, and to understand if there might be versions of the same symmetry in general CFTs. Is there a way to construct an object corresponding to the spin-shadow of a local operator353535or, more precisely, in the framework of Kravchuk:2018htv one should investigate the shadow transform applied to the light transform of a local operator. proving that it should sit on the Chew-Frautschi plot?
@@
An interesting problem is also raised by the increase of degeneracies of trajectories with twist. What does this imply together with the symmetry? Two possible mechanisms likely both play a role: first, some higher-twist trajectories could have points below the unitarity bound that, despite having negative integer spin (in steps of 2), are not spin-flipped images of local operators, and, from our point of view, this would be highlighted by the fact that they satisfy QSC equations with non-constant gluing matrix; [^36] We can expect that such points would also have exact zeros for the analytic continuation of the OPE coefficients. This comes from extending below unitarity the principle that a non-constant gluing matrix is equivalent to the divergence of the norm $\langle\mathbb{O}\mathbb{O}\rangle$ of Henriksson:2023cnh , which then implies the vanishing of the OPE coefficient. In Fig. 2 of Julius:2024ewf we see already some examples of this mechanism: in the case of the four subleading trajectories studied there, we see that points where the OPE coefficients vanish are precisely the ones that are neither local operators nor their spin-shadows below unitarity, i.e. they are not involved in the twist/co-twist symmetry. in addition, some “extra” trajectories might bend before reaching these negative values of $S$. In general, it would be important to decode the full pattern in which trajectories are threaded together by the twist/co-twist symmetry.
@@
Finally, another motivation for developing tools for studying continuous spin is to gather a new type of data to use in the Bootstrability approach Cavaglia:2021bnz ; Cavaglia:2022qpg ; Cavaglia:2023mmu ; Caron-Huot:2022sdy . So far, this program has used the spectrum of low-lying local operators coming from QSC methods as an input for the conformal bootstrap in order to derive constraints on correlation functions. Since continuous spin comes with a tight structure in CFT Caron-Huot:2017vep , it would be very nice to explore if the input on Regge trajectories at any continuous $S$ coming from the QSC could be used to put even more constraints on OPE coefficients. This direction might be important especially to study Bootstrability for the full 3D or 4D theory (see Caron-Huot:2022sdy ), which is significantly more complex than the most studied defect Wilson line setup. In a parallel development, in the last few years there was significant progress in the study of 4-point correlators in $\mathcal{N}$$=$$4$ SYM at strong coupling Alday:2022uxp ; Alday:2022xwz ; Alday:2023mvu ; Alday:2023jdk ; Alday:2023flc . The approach of these works uses the analytic conformal bootstrap, with input from integrability spectral data on the first Regge trajectories. This analysis has also led to new structural insight on the analytic structure of the spectrum itself Julius:2023hre ; Julius:2024ewf . The symmetry we are discussing might give further constraints and be helpful in this program, as well as for the one recently initiated for ABJM theory in Chester:2024esn .
@@ Acknowledgements
We are grateful to Mikhail Alfimov, Francesco Aprile, Till Bargheer, Benjamin Basso, Davide Bonomi, Carlos Bercini, Lorenzo Bianchi, Riccardo Conti, Nikolay Gromov, Tobias Hansen, Johan Henriksson, Julius Julius, Rob Klabbers, Shota Komatsu, Petr Kravchuk, Madalena Lemos, Marco Meineri, Michelangelo Preti, Nicolò Primi, Nika Sergeevna Sokolova, Istvan Szećsényi for discussions related to this work.
@@ Appendix A Weak coupling checks of the symmetry in ABJM theory
$\gamma_{2}$ $=4\left(S_{1}(j+1)+S_{-1}(j+1)\right)+\frac{2\left(1-(-1)^{1+j}\right)}{j+2}+8,$ $\text{for}\,\,\mathcal{V}^{4}_{3+2j,3},\quad(\tau=2,\,j\in\mathbb{N}),$
$\gamma_{2}$ $=4\left(S_{1}(j+2)-S_{-1}(j+2)\right)+\frac{\left(1+(-1)^{j}\right)}{j+3}-\frac{\left(1-(-1)^{j}\right)}{j+2}+4,$ $\text{for}\,\,\mathcal{V}^{4}_{4+2j,4},\quad(\tau=3,\,j\in\mathbb{N}),$
$\gamma_{2}$ $=4\left(S_{1}(j+2)+S_{-1}(j+2)\right)+\frac{2\left(1-(-1)^{j}\right)}{j+3}+\frac{32}{3},$ $\mathcal{V}^{4}_{5+2j,5},\quad(\tau=4,\,j\in\mathbb{N}_{\text{odd}}),$ (194)
@@
$j=\frac{\text{max}\left(\delta_{1},0\right)-\text{max}\left(\delta_{2},0\right)}{2},\quad\Delta=L+\frac{\text{max}\left(\delta_{1},0\right)+\text{max}\left(\delta_{2},0\right)}{2}.$ (196)
@@
$\gamma_{2}=4\left(S_{1}(j+1)-S_{-1}(j+1)\right),\quad\text{for\,\,}\mathcal{V}^{2}_{3+2j},\quad(\tau=1,\,j\in\mathbb{N}).$
@@
$\left.\mathcal{V}^{2}_{3+2j}\right|_{j\rightarrow 0},\quad(\tau=1,\tilde{\tau}=2)\quad\leftrightarrow\quad\left.\mathcal{V}^{4}_{3+2j,3}\right|_{j\rightarrow-1},\quad(\tau=2,\tilde{\tau}=1).$ (197)
@@
$\left.\mathcal{V}^{4}_{3+2j,3}\right|_{j\rightarrow 0},\quad(\tau=2,\tilde{\tau}=3)\quad\leftrightarrow\quad\left.\mathcal{V}^{4}_{4+2j,4}\right|_{j\rightarrow-1},\quad(\tau=2,\tilde{\tau}=2),$ (198)
@@
$\left.\mathcal{V}^{4}_{4+2j,4}\right|_{j\rightarrow 0},\quad(\tau=3,\tilde{\tau}=4)\quad\leftrightarrow\quad\left.\mathcal{V}^{4}_{5+2j,5}\right|_{j\rightarrow-1},\quad(\tau=4,\tilde{\tau}=3),$ (199)
@@
We also extracted numerically the two-loop anomalous dimension of the point $(\tau=5,\tilde{\tau}=2)$ living, below the unitarity bound, on the subleading trajectory shown in Figure 12 at $S_{sl(2)}=-1$. This gives an excellent agreement with $\gamma_{2}=12$, which is the two-loop anomalous dimension of the physical operator with $S_{sl(2)}=2$ ($\tau=2,\tilde{\tau}=5$) belonging to the $\mathcal{V}^{4}_{3+2j,3}\big{|}_{j\to 1}$ multiplet Papathanasiou:2009zm .
@@
Notice that in general the symmetry connects a state of spin $j$ on a trajectory of twist $\tau$ to a point on a trajectory of twist $\tilde{\tau}=d-2+\tau+2j$, continued to unphysical spin $\tilde{j}=-d+2-j$. The only instances of the symmetry which we can check based on the analytic trajectories (A) are the ones we have presented, as we do not have the exact expressions in order to perform the analytic continuations to negative $j$ for higher twists. However, it would be interesting to use numerical explorations of the QSC solutions to map the precise pairings of multiplets given by the symmetry.
@@ Appendix B Useful ancillaria
B.1 Leading coefficients
Here we report some additional internal constraints of the QSC construction. The leading prefactors coefficients of the expansion (174) satisfy (see Bombardelli:2017vhk )
@@ B.1 Leading coefficients
Up to Weyl reflections 2.2, the only points satisfying the BPS constraint (190) are
$(\Delta_{sl(2)},S_{sl(2)})\in\{(1,0),(0,-1)\}.$ (201)
Studying the Q-system equations at large $u$ one can constrain the form of the coefficients $B_{a|i}$ entering in the expansion (18). The expressions are provided below in Mathematica format, where $B_{a|i}\equiv\texttt{B}\texttt{[}a,i\texttt{]}$, the charges are denoted as $M_{A=1,2,5}\equiv\texttt{M}\texttt{[}A\texttt{]}$ and $\hat{M}_{I=1,2,5}\equiv\texttt{Mh}\texttt{[}I\texttt{]}$, and the coefficients $\mathcal{A}_{ab}\equiv\texttt{A}\texttt{[}a,b\texttt{]}$.
{B[2, 1] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[2, 4]
B[1, 1]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[2, 2] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[2, 4]
B[1, 2]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] - Mh[2]) (M[5] - Mh[2])),
B[2, 3] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[2, 4]
B[1, 3]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[2, 4] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[2, 4]
B[1, 4]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] - Mh[2]) (-M[5] + Mh[2])),
B[3, 1] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[3, 4]
B[1, 1]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[3, 2] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[3, 4]
B[1, 2]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] - Mh[2]) (M[5] - Mh[2])),
B[3, 3] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[3, 4]
B[1, 3]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[3, 4] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 3] A[3, 4]
B[1, 4]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] - Mh[2]) (-M[5] + Mh[2])),
B[4, 1] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 4] A[3, 4]
B[1, 1]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[4, 2] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 4] A[3, 4]
B[1, 2]) / ((M[2] + Mh[1]) (M[5] + Mh[1]) (M[2] - Mh[2]) (M[5] - Mh[2])),
B[4, 3] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 4] A[3, 4]
B[1, 3]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] + Mh[2]) (M[5] + Mh[2])),
B[4, 4] -> ((M[1] + M[2]) (M[2] - M[5]) (M[1] + M[5]) (M[2] + M[5]) A[2, 4] A[3, 4]
B[1, 4]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] - Mh[2]) (-M[5] + $\quad\texttt{B[1, 4]) / ((M[2] - Mh[1]) (-M[5] + Mh[1]) (M[2] - Mh[2]) (-M[5] + Mh[2]))}\texttt{\big{\}}},$
together with
B[1, 4] -> ((M[1] - Mh[1])^2 (M[2] - Mh[1])^2 (M[5] - Mh[1])^2 (M[1] + Mh[1])^2)
(M[2] + Mh[1])^2 (M[5] + Mh[1])^2 (M[1] - Mh[2])^2 (M[2] - Mh[2])^2)
(M[5] - Mh[2])^2(M[1] + Mh[2])^2 (M[2] + Mh[2])^2 (M[5] + Mh[2])^2)/
(16 (M[1] - M[2])^2(M[1] + M[2])^4 (M[1] - M[5])^2(M[2] - M[5])^2)
(M[1] + M[5])^4 (M[2] + M[5])^4 Mh[1]^2 Mh[2]^2(Mh[1]^2 - Mh[2]^2)
[…]
B.2 Generic form of the Weyl reflections
For completeness, we report a slightly more general form of the Q-system transformations implementing the Weyl reflections, which simply differs from the one given in the main text by a generic unphysical “gauge transformation”, i.e. a redefinition of the normalisations of various Q-functions. In general, the transformation is implemented as
$\mathbf{Q}\rightarrow\mathbf{Q}^{\prime}\equiv\mathbb{M}\cdot\mathbf{Q}\cdot\mathbb{M}^{T},$ (202)
and correspondingly
$Q_{a|i}\rightarrow Q^{\prime}_{a|j}=Q_{a|j}\mathbb{M}_{i}^{j}\,.$ (203)
The most general form of $\mathbb{M}$ implementing the Weyl transformation $\texttt{W}_{S}$ or $\texttt{W}_{\Delta}$ takes the form, respectively,
$\left(\mathbb{M}_{1}\right)_{i}^{\,j}=\begin{pmatrix}a_{1}&0&0&0\\ 0&0&a_{2}&0\\ 0&-\frac{1}{a_{2}}&0&0\\ 0&0&0&\frac{1}{a_{1}}\end{pmatrix}\,\leftrightarrow\,\texttt{W}_{S},$ (204)
$\left(\mathbb{M}_{2}\right)_{i}^{\,j}=\begin{pmatrix}0&0&0&b_{1}\\ 0&b_{2}&0&0\\ 0&0&\frac{1}{b_{2}}&0\\ -\frac{1}{b_{1}}&0&0&0\end{pmatrix}\,\leftrightarrow\,\texttt{W}_{\Delta},$ (205)
where the matrices given in (49) are simply a particular gauge choice. In particular, with the transformation $\mathbb{M}_{1}$ in (204), on top of the interchange of columns with $i=1$ and $i=4$, the prefactors $B_{a|i}$ in the normalisation of Q-functions are redefined according to
$a_{1}=\frac{\left(B_{a|1}\right)^{\prime}}{\left(B_{a|1}\right)}=\frac{\left(B_{a|4}\right)}{\left(B_{a|4}\right)^{\prime}},\quad a_{2}=\frac{\left(B_{a|2}\right)^{\prime}}{\left(B_{a|3}\right)}=-\frac{\left(B_{a|2}\right)}{\left(B_{a|3}\right)^{\prime}}.$ (206)
Similarly, the transformation $\mathbb{M}_{2}$ interchanges columns $i=2$ and $i=3$ with the normalisations redefined as
$b_{1}=\frac{\left(B_{a|1}\right)^{\prime}}{\left(B_{a|4}\right)}=\frac{\left(B_{a|1}\right)}{\left(B_{a|4}\right)^{\prime}},\quad b_{2}=\frac{\left(B_{a|2}\right)^{\prime}}{\left(B_{a|2}\right)}=-\frac{\left(B_{a|3}\right)}{\left(B_{a|3}\right)^{\prime}}.$ (207)
The redefinition of the prefactors of $\mathbf{Q}_{ij}$ functions also follows easily.